0.1 ml of a 10-4 dilution of M. luteus was plated. After incubation, 80 colonies were counted. How many colony-forming units per ml were in the original E. coli sample?
Answer #1

CFU per mL = number of colonies / (volume plated × dilution)

Colonies = 80

Dilution = 10^-4

Volume plated = 0.1 mL

CFU per mL = 80/(10^-4 × 0.1) = 80 × 10^5 = 8 × 10^6 CFU per mL

Answer is 8 × 10^6 CFU per mL

Please rate high.